Design Issues in Using the WWW for Education

Bonnie Mitchell

Syracuse University

Abstract

This presentation will cover the important concepts and technical aspects of using the WWW for education. Issues such as interface design, expressive content, and technical development will be illustrated. General concepts discussed: hypermedia, intuitive vs exploratory interfaces, elements and principles of aesthetic interface design, expressing content, engaging interest, involving meaningful interaction.

Coffee and cookies will be served